the perimeter of a triangle is 480 meters. the second side is three times the first and the third side is sixty meters more than the second. what are the lengths of each side

Answer by checkley79(3341)

P=A+B+C
B=3A
C=B+60 OR C=3A+60
480=A+3A+3A+60
480=7A+60
7A=480-60
7A=420
A=420/7
A=60 ANS.
B=3*60=180 ANS.
C=3*60+60=180+60=240 ANS.
PROOF:
60+180+240=480
480=480